Consider the function (x) = x - 2x-5 What is (-1)? What is r(5) ? What is limy 0+ F(x)? Enter a number.
Answer #1

Given f(x) lxl- /2x-51 x- 5 when no -) fox) - when 3x - 5 osn < 2 -xt 5 when ant Now Hel=0 - 705) # = um x ot 3